Interest centers around the life of an electronic component. Suppose it is known that the probability that the component survives for more than 7000 hours is 0.49

Suppose also that the probability that the component survives no longer than 2000 hours is 0.04

​(a) What is the probability that the life of the component is less than or equal to 7000 ​hours?

​(b) What is the probability that the life is greater than 2000 hours?
Answer #1

a) P(life of the component is less than or equal to 7000 ​hours) = 1 - P(component survives for more than 7000 hours)

= 1 - 0.49

= 0.51

b) P(life is greater than 2000 hours) = 1 - P(the component survives no longer than 2000 hours)

= 1 - 0.04

= 0.96